# General Riemann Integration on Partially Ordered Net Convergence Structures

### v1

@@ 5 Pavlakos Lebesgue Type Integration and its Riemmanian Properties
We now expound some aspects of a very specific type of Lebesgue integral, by which we mean integrals defined by approximating the functions to be integrate by a net of simple/elementary functions, for functions and measures with values in Riesz spaces. For a brief and incomplete survey of such constructions, we mention the uniform net limit type integrals of [81], [30], [31], [29], the sequential integrals of [32], [82], [83], [84], [85], [86], [87], [88], [89], [90], [91], [92], [93], [94], [95], [96], [97], [98] [38], [99], [100], [57], [12], [39], [101], [102], [103], [104], [105], [42], [43], [41] the theses of [106], [107], [108], [109], containing abundant information on the structures that will be discussed here, as well as the related Daniell type integrals, not covered here in any detail, of [33], [52], [110], [111], [112], [113], [114], [115], [116], [117], [118], [119], [120], [121], [122], [123], [124], [125], [126], [127], [128], [129], [130], and [131].
@@
    Riesz spaces $X,Y,Z$ such that $Z$ is Dedekind complete and weakly $\sigma$-distributive. In relation to the measure theoretic objects, $(\Omega,\mathcal{H})$ will always denote a measurable space, i.e, $\mathcal{H}$ is a $\sigma$-algebra of subsets of $\Omega$, and $\mu:\mathcal{H}\rightarrow Y^{+}$ will be $\sigma$-additive in the net convergence structure of $Y$ in the sense of Definition 3.13. We also suppose that $\mu$ is a regular integrator.
@@
With these observations, we pass to a brief presentation the integration theory of Pavlakos [30], which contains the integral of Kappos [29] and Berg [81].
To develop his integral, Pavlakos [30] supposes that $X$ is of countable type with respect to order convergence, i.e he supposes $X$ is fist countable with respect to the net convergence structure given by order convergence. Therefore, we will also suppose throughout this section, besides 1. and 2. above, that:
@@
Now, we define a list of functions spaces specified by Pavlakos in different sections in [30]. To obtain clarity and brevity, we will define the function spaces in question already with the hypothesis put onto them by Pavlakos [30] for the purpose of integration, and with the same notation.
@@
    $F(\Omega,X)$, the space of all functions $f:\Omega\rightarrow X$ equipped with addition and multiplication by (real) scalars pointwise, and with the pointwise partial order inherited from $X$. In this case, it becomes a Riesz Space.
@@
    uniformly almost everywhere in order as defined in the previous section, or $(u)-\lim_{n\rightarrow\infty}f_{n}=f$ a.e.-$\mu$ on $\Omega$ in [30] notation and terminology. This is called the space of $(\mathcal{H},\mu)$-measurable functions.
@@
with the same notation as in the previous definition, we say that it is Pavlakos integrable if the following (unconditional series!) limit:
@@
By Lemmas 1.2 and 3.1 of [30], this quantity is well defined in the sense of not depending on the rearrangement of the series and the representation of $f$ as an elementary function (which is, in general, non-unique).
@@
By Theorem 3.5 and Corollary 3.7 of [30], the defined integral on $E(\Omega,T)$ is a linear, positive and isotone operator and it is absolute, in the sense that $f$ is Pavlakos integrable if and only if $|f|$ is Pavlakos integrable.
We now enounce Theorems 3.11 and Corollary 3.12 of [30], respectively, for the definition of the integral on $M(\Omega,X)$:
@@
We then have the main integral of [30]:
@@
It is a consequence of Theorem 5.1 and Lemma 5.1 that this integral is well defined. The definition of this integral on a set $A\in\mathcal{H}$ and its properties are as same as before. For more basic properties, we refer the reader to the original article of Pavlakos [30].
The last special property of this integral that we will need is the following Theorem 3.8 of [30]:
@@
is $\sigma$-additive in $\mathcal{H}$.
@@
where $b_{i}$ is the value of $f$ at $B_{i}$.
@@
$\int_{\Omega}^{S^{*}}fd\mu=\sum_{n=1}^{\infty}a_{n}\mu(A_{n})$
@@
For the general case, we simply write the decomposition of a Pavlakos integrable $f$ on its negative and positive parts, both non-negative and Pavlakos integral, use Theorem 5.3, and then we get:
@@
One interesting aspect of the discussion from now on is, then, to connect net-approximation (Lebesgue) type integrals to the Pavlakos’ one in an attempt to relate them to the Riemmanian approach. As real measurable non-negative functions $f:\Omega\rightarrow\mathbb{R}^{+}$ are strongly measurable in the sense of Böchner (see [132]), and thus can be approximated uniformly by a sequence of increasing elementary functions (see again [132]), we may use the monotone convergence theorem of Maitland-Wright [32] and the device of decomposing a function into negative and positive parts to prove:
@@
For more details on the Maitland-Wright integral, see [32] and [82], and [30]. A similar result can be obtained for the integral of Kusraev and Tasoev [102].
This result also helps to connect the Riemmanian approaches to integration more closely with the sequential/net-based Lebesgue integrals in the most well known approaches to integration in partially ordered spaces, at least for real functions. We now pass to a brief discussion of a general case for Lebesgue and Saks-type integration, and comment more about the still-to-be established results classifying all these integration procedures.
6 General Lebesgue and Saks Type Integration on Boccuto-Candeloro Net Convergence Structures
6.1 Abstract Lebesgue Integration
Taking the cues in the last section, we see that, in general, all the Lebesgue type procedures for integration cited previously follows the same methodology:
@@ 6.1 Abstract Lebesgue Integration
    Choose a space of functions $L$ that contains the simple (measurable) functions and define the integral for this former class in some sense. For the simple functions we define the integral, generally, as a linear combination of the constants and measure of the sets in the definition.

### v2

@@ Abstract
We now expound some aspects of a very specific type of Lebesgue integral for functions and measures with values in Riesz spaces. For a brief and incomplete survey of such constructions, we mention the uniform net limit type integrals of Berg (1978), Pavlakos (1983), Pavlakos (1991), Kappos (1970), the sequential integrals of Wright (1969), Maitland-Wright (1972), Sipos (1981), Potocky (1972), Potocky (1977), Potocký (1986), Maličký (1986), Maličký (1987), Maličký (1990), Sobolev and Shcherbin (1975), Sobolev and Shcherbin (1992), Cristescu (1976), Cristescu (1994), Cristescu (2001), Aleksyuk and Isakov (1982), Poroshkin and Bazhenov (1982), Haluška (1993), Andreu et al. (1989) Boccuto (1995), Boccuto (1997), Boccuto (1998), Boccuto and Candeloro (2009b), Boccuto and Dimitriou (2015), Boccuto and Sambucini (2023), Kusraev and Malyugin (1988), Kusraev and Tasoev (2017), Grobler and Labuschagne (2017), van Rooij and van Zuijlen (2016), van Rooij and van Zuijlen (2017), de Jeu and Jiang (2022a), de Jeu and Jiang (2022b), Jiang et al. (2024) the theses of Pavlakos (1972), Jeurnink (1982), Groenewegen (1982), van Zuijlen (2012), containing abundant information on the structures that will be discussed here, as well as the related Daniell type integrals, not covered here in any detail, of McShane (1953), Matthes (1961), Kappos and Mallios (1960), Mallios (1960), Sisto (1993), Leontiadis (1986), Leontiadis (1989), Leontiadis (1994), Alfsen (1958), Higgs (1980), Wilhelm (1984), Wilhelm (1988), an der Heiden (1970), Shamaev (1986), Vonkomerova (1981), Congost Iglesias (1980), Riecan (1975), Riecan (1982), Riecan (1983), Grobler (2021), Vrábel and Vrábelová (2011), Isakov (2005), Markwardt (1986), and Westerbaan (2019).
@@
    Riesz spaces $X,Y,Z$ such that $Z$ is Dedekind complete and weakly $\sigma$-distributive. In relation to the measure theoretic objects, $(\Omega,\mathcal{H})$ will always denote a measurable space, i.e, $\mathcal{H}$ is a $\sigma$-algebra of subsets of $\Omega$, and $\mu:\mathcal{H}\rightarrow Y^{+}$ will be $\sigma$-additive in the (partially ordered) net convergence structure of $Y$ in the sense of Definition 3.4. We also suppose that $\mu$ is a regular integrator.
@@
With these observations, we pass to a brief presentation the integration theory of Pavlakos Pavlakos (1983), which contains the integral of Kappos (1970) and Berg (1978).
To develop his integral, Pavlakos (1983) supposes that $X$ is of countable type with respect to order convergence, i.e $X$ is fist countable with respect to the net convergence structure given by order convergence (see O’Brien et al. (2023)). Therefore, we will also suppose throughout this (sub)section, besides 1. and 2. above, that:
@@
Now, we define a list of functions spaces specified by Pavlakos in different sections in Pavlakos (1983). To obtain clarity and brevity, we will define the function spaces in question already with the hypothesis put onto them by Pavlakos (1983) for the purpose of integration, and with the same notation.
@@
    $F(\Omega,X)$, the space of all functions $f:\Omega\rightarrow X$ equipped with addition and multiplication by (real) scalars pointwise, and with the pointwise partial order inherited from $X$. In this case, this function space becomes a Riesz Space.
@@
    uniformly almost everywhere in order as defined in the previous section, or $(u)-\lim_{n\rightarrow\infty}f_{n}=f$ a.e.-$\mu$ on $\Omega$ in Pavlakos (1983) notation and terminology. This is called the space of $(\mathcal{H},\mu)$-measurable functions.
@@
with the same notation as in the previous definition, we say that it is Pavlakos integrable if the following (unconditional series) limit:
@@
By Lemmas 1.2 and 3.1 of Pavlakos (1983), this quantity is well defined in the sense of not depending on the rearrangement of the series and the representation of $f$ as an elementary function (which is, in general, non-unique).
@@
By Theorem 3.5 and Corollary 3.7 of Pavlakos (1983), the defined integral on $E(\Omega,T)$ is a linear, positive and isotone operator and it is absolute, in the sense that $f$ is Pavlakos integrable if and only if $|f|$ is Pavlakos integrable.
We now enounce Theorems 3.11 and Corollary 3.12 of Pavlakos (1983), respectively, for the definition of the integral on $M(\Omega,X)$:
@@
We then have the main integral of Pavlakos (1983):
@@
It is a consequence of Theorem 5.1 and Lemma 5.1 that this integral is well defined. The definition of this integral on a set $A\in\mathcal{H}$ and its properties are as same as before. For more basic properties, we refer the reader to the original article of Pavlakos (1983).
The last special property of this integral that we will need is the following Theorem 3.8 of Pavlakos (1983):
@@
is $\sigma$-additive on $\mathcal{H}$.
@@
where $b_{i}$ is the value of $f$ at $B_{i}$, where all the series are unconditionally convergent.
@@
$\int_{\Omega}^{S^{*}}fd\mu=\sum_{n=1}^{\infty}a_{n}\mu(A_{n}),$
where all the series are unconditionally convergent.
@@
For the general case, we simply write the decomposition of a Pavlakos integrable $f$ on its negative and positive parts, both non-negative and Pavlakos integrable, use Theorem 5.3, and then we get:
@@
One interesting aspect of the discussion from now on is, then, to connect net-approximation (Lebesgue) type integrals to the Pavlakos’ one in an attempt to relate them to the Riemmanian approach. As real (Borel) measurable functions $f:\Omega\rightarrow\mathbb{R}$ are strongly measurable in the sense of Böchner (see Dinculeanu (2000)), and thus can be approximated uniformly by a sequence of increasing elementary functions (see again Dinculeanu (2000) and monteirodissertação for a study of how uniform convergence in the usual sense, for real functions, is equivalent to ours), we may use the monotone convergence theorem of Maitland-Wright proved in Wright (1969) and the device of decomposing a function into negative and positive parts to prove:
Proposition 5.2 (Maitland-Wright Integrable Function is Pavlakos Integrable).
As a direct corollary, we have:
Proposition 5.3 (Maitland-Wright Integrable Function is $S^{*}$ Integrable).
Let $f:\Omega\rightarrow\mathbb{R}$ be a Maitland-Wright integrable function on $\Omega$, then it is $S^{*}$ integrable on $\Omega$ and the two integrals agree.
For more details on the Maitland-Wright integral, see Wright (1969) and Maitland-Wright (1972), and Pavlakos (1983). A similar result can be obtained for the integral of Kusraev and Tasoev constructed in Kusraev and Tasoev (2017).
Thus, the results above helps to connect the Riemmanian approaches to integration more closely with the sequential/net-based Lebesgue integrals in the most well known approaches to integration in partially ordered spaces, specially for real functions and Riesz-valued measures and Riesz-valued functions and measures with uniform approximations. A further interesting aspect would be to make analogous Riemmanian constructions of the integral of de Jeu and Jiang (2022b), which can take values in the Dedekind completion of a given Riesz space.
5.2 Lebesgue Integrals in Topological Vector Spaces and Riemmanian Equivalences.
On this section, we will study some aspects of the general Lebesgue integral, previously defined, but now in the setting of, not necessarily partially ordered, topological vector spaces. The literature on the subject contains numerous publications, most of them based on the classical integrals of Birkhoff (1935), Bochner (1933) and Pettis (1938). A great bibliography in this context is provided in the book Henstock (1991).
Nevertheless, there is a less know part of the subject which begins with the construction of the remarkable bilinear integrals of Dobrakov (1970) and Bartle (1956) (see also Dobrakov and Panchapagesan (2004), Panchapagesan (1995) and Panchapagesan (2008)), and contains reverberations in the recent literature, specially in view of applications in mathematical physics (see Jefferies (2017)). The period after the articles of Dobrakov and Bartle just cited contained different generalizations, and in different directions (again, non-exhaustively, but not covered by Henstock (1991)):
  1. (i)
    For (real) scalar functions integrated with respect to (convexly bounded) vector measures with values in a topological vector space, we have the recent works of Labuda (2010), Labuda (2014) and Drewnowski and Labuda (2013), which generalized the earlier defined integrals by Turpin (1975), Rolewicz (1985), Butković (1979), Butkovic (1979) and Thomas (2012) (based on Thomas (1972)).
  2. (ii)
    A student of Dobrakov, Jan Haluska, beginning in Haluška (1988) and Haluška (1997), developed a generalization of the Dobrakov procedure for locally convex bornological topological vector spaces. This integral contains structures of product type and various convergence theorems, which can be gauged from Haluška and Hutník (2010) and Haluška and Hutník (2012). Connections with the $S^{*}$-partition integral were made in Duchoň and Haluška (2005), which will be commented below. Related integrals are considered in Dunford (1975), Sivasankara (1981), Chakraborty and Basu (2011), Chakraborty and Basu (2009), Chakraborty and Basu (2007), Roy and Chakraborty (1986), Rao Chivukula and Sastry (1983), Debieve (1981) and Kappeler (1981).
  3. (iii)
    Related bilinear vector integrals, specially based on tensor products, were developed by B. Jefferies, S. Okada, B. Ricker and I. Kluvanek, whose works are summarized in the books Jefferies (1996), Kluvánek (1988) and Jefferies (2017). An important contribution of the integration procedure of these authors is to have a smaller space of integrable functions, but without supposing finite semivariation of the (operator valued) measures involved. This has important applications in mathematical physics as emphasized by Jefferies (2017).
  4. (iv)
    The "integration school" of Bombal, Rodriguez-Salinas and Jimenez-Guerra in Spain developed similar procedures in bornological (locally convex) spaces with the publications of Bombal Gordon (1981a), Bombal Gordon (1981b), Rodríguez Salazar (1985), Rodríguez-Salinas (1983a), Rodríguez-Salinas (1983b), Rodríguez-Salinas (1983c), Rodríguez-Salinas (1983d), Ballvé (1988), Ballvé and Jiménez Guerra (1993), José Fernandez y Fernandez-Arroyo (1987), Fernández y Fernández-Arroyo (1992), Fernández y Fernández-Arroyo et al. (1995), Bravo de La Parra (1987), Hidalgo Alonso (1994). Basically, this is a type of integral in locally convex spaces which satisfy an Egorov approximation procedure.
  5. (v)
    With a view towards stochastic integration, Brooks and Dinculeanu (1974) developed a bilinear integration theory based in a family of Banach-valued set functions, possibly finitely additive, which lead Dinculeanu (2000) to a very general and applicable integration theory. A more recent development is included in the thesis of Kozinski (2005).
  6. (vi)
    Following the ideas of Drewnowski (1972b) in integration theory, W. Marik (in Germany) and J-C. Massé (in Canada) independently developed a bilinear integration theory for semigroup-valued measures and functions based on convergence in (a family of) submeasure(s). The initial developments are contained in the works Marik (1979) and Massé (1974), and later in a reduced version in Marik (1987) and Masse (1979). A special case of this theory is the integration of Heinich (1971) and Ohba (1982), already cited above, in groups based on uniform approximations. A very interesting development of this theory for conoids is given in the thesis of Abreu (2006), and in Banach spaces by Agafonova (1976). An independent axiomatic development of these types of submeasure based integrals with values in semigroups has also been developed by Isakov (2005), as well as in the unpublished integration material in Drewnowski (1971).
  7. (vii)
    Based on initial ideas of Schäfke (1970) and Schäfke (1971), a very general integration procedure for lattice normed spaces, and even partially ordered semigroup-valued measures without any restriction on the semivariations, have been developed by Schäfke (1972a), Schäfke (1972b), Schaefke (1977), Hoffmann and Schäfke (1992), with ramifications in absolute Riemann integrability in Günzler (1985), Guenzler (1980), Günzler (2001), Díaz Carrillo and Günzler (1993), Díaz Carrillo (2002), Diaz Carrillo and Muñoz Rivas (1994). The most general form of these results are contained in the works of Volkmer and Weber (1982), Weber (1977a), Weber (1977b), Markwardt (1986), Markwardt (1981) and Hoffmann (1977).
In view of this great variety, it would take a more voluminously work to deal with all these integration procedures. We will privilege those that are more related to Riemmanian procedures constructed in the previous sections, and in order of complexity we shall construct the following scheme:
  • 1.
    First, we will show that the Bartle-Dunford-Schwartz of Drewnowski and Labuda (2013) is contained in the net Riemann integral in the form of the $S^{*}$-partition integral in (sequentially complete) Hausdorff topological vector spaces.
  • 2.
    Then, we will briefly expose the results of Duchoň and Haluška (2005), which shows that the bornological integral of Haluska, and therefore the special case of Dobrakov Banach-valued integral, is a special case of the $S^{*}$-partition integral in bornological locally convex topological vector spaces.
  • 3.
    Lastly, we construct some elements of the integration theory of Masse (1979), showing it contains the related integral of Marik (1987), as well as the previous Lebesgue ones cited so far in this list, and prove partial results of the connection of this integration procedure with the net Riemann integral.
To not increase the technicalities involved, we will deal with the case of topological vector spaces, leaving (uniform) semigroups for a further work. Also, most of these theories have convergence theorems of dominated and Vitali type, but unless necessary, we will not formulate them explicitly. For every non-defined term concerning topological vector spaces, bornologies and convergences in them, see Beer (2023), Schaefer and Wolff (1999), Radyno (1982), Jarchow (1981) and Hogbe-Nlend (1977).
5.2.1 The Bartle-Dunford-Schwartz (BDS) in Topological Vector Spaces.
In this section, we follow mainly the ideas and concepts of Labuda (2014) and Drewnowski and Labuda (2013).
We fix, for this whole section, $(X,\tau)$ a Hausdorff topological vector space with a topology $\tau$ that is sequentially complete and has $\mathcal{U}$ any base of balanced zero-neighborhoods in $X$ and $Fs(X)$ any base of continuous $F$-seminorms which define its topology. Fix also $\mathcal{H}$ a $\sigma$-algebra of subsets of a (non-empty) set $\Omega$.
The main definition for measures we shall use is the following:
Definition 5.4 (Convexly Bounded Vector Measures).
We say that a vector measure (that is, a $\sigma$-additive set function) $\mu:\mathcal{H}\rightarrow X$ is convexly bounded if the convex hull of its image is bounded [^20] In the bornological, or Von Neumann, sense..
An equivalent definition which we shall use employs the continuity properties of simple-integral operators:
@@
A vector measure $\mu:\mathcal{H}\rightarrow X$ is convexly bounded if and only if the operator:
$\int_{\Omega}\cdot\ d\mu:(S_{\infty},||\cdot||_{\infty})\rightarrow(X,\tau),$
given by:
$\int_{\Omega}sd\mu=\sum_{i=1}^{n}s_{i}\mu(A_{i}),$
is continuous, where $(S_{\infty},||\cdot||_{\infty})$ is the (Banach) space of simple functions equipped with the supremum norm.
The proof follows directly from the relation between continuity of operators and bornologies between two topological vector spaces.
In this case, we get the connection of finite semivariation defined previously:
Proposition 5.5 (Convex Boundedness Implies Finite Semivariation).
If a vector measure $\mu:\mathcal{H}\rightarrow X$ is convexly bounded, then it has finite semivariation in the sense of Definition 4.7.
With these properties in mind, we now define a sequence of objects in view of the BDS integral of Labuda (1976). For now on in this section, $\mu$ will be a convexly bounded vector measure.
In this case, let, for $d(\cdot)$ an $F$-seminorm in $Fs(X)$ and $A\in\mathcal{H}$,
$d(\mu)(A)=\sup\{d(\mu(B)):B\in\mathcal{H},B\subseteq A\},$
the submeasure-majorant associated to $\mu$ and $d$. It’s readily seen that the submeasure-majorant is an order continuous (real) $\sigma$-subadditive set function defined on $\mathcal{H}$.
We then say that a real function $f:\Omega\rightarrow\mathbb{R}$ is $\mu$-measurable if it is equal, except on a $\mu$-null set [^21] That is, an $A\in\mathcal{H}$ such that $\mu(B)=0$ in $Y$ for every $B\subseteq A$ such that $B\in\mathcal{H}$., to a (Borel) $\mathcal{H}$-measurable function $g$. We denote the (quotient space by $\mu$-null sets of the) space of all such $\mu$-measurable functions by $L^{0}(\mu)$. We equip this space with the Hausdorff vector topology of convergence in (sub)measure $\mu$ (see Labuda (2014) for details).
In this case, we now have the main definition from Drewnowski and Labuda (2013):
Definition 5.5 (Bartle-Dunford-Schwartz Integral).
A function $f\in L^{0}(\mu)$ is said to be Bartle-Dunford-Schwartz (BDS) integrable if there exists a sequence $\{f_{n}\}_{n\in\mathbb{N}}$ of simple functions such that $f_{n}\rightarrow f$ $\mu$-a.e. and, for each $A\in\mathcal{H}$, $\lim_{n\rightarrow\infty}\int_{A}fd\mu$ exists in $X$, where each $\int_{A}f_{n}d\mu$ is defined additively.
Then, by definition,
$\int^{(BDS)}_{A}fd\mu=\lim_{n\rightarrow\infty}\int_{A}fd\mu.$
This integral has the usual additivity properties in the space of integrable functions, and also good continuity properties (see sections $3$ and $4$ of Drewnowski and Labuda (2013)).
Observation 5.1 (The BDS integral Ignores $\mu$-null sets).
By the definition of the $BDS$-integral, we have that, if $A\in\mathcal{H}$ is a null set, then for every $f\in L^{0}(\mu)$,
$\int_{A}^{(BDS)}fd\mu=\int_{\Omega}^{(BDS)}f\mathbbm{1}_{A}d\mu=0.$
In this case, if $f$ is $BDS$-integrable and $A$ is the $\mu$-null set on which the convergence of $\{f_{n}\}_{n\in\mathbb{N}}$ occurs outside it, we have (see also more details in Drewnowski and Labuda (2013)):
$\int_{B}^{(BDS)}fd\mu=\int_{B}^{(BDS)}f\mathbbm{1}_{A^{c}}d\mu,$
for every $B\in\mathcal{H}$. As the same property of integration on null sets is valid for the $S^{*}$-integral by the results already stated and used in Section 4.1, we may suppose without loss of generality that $f$ is, itself, Borel-measurable to prove the equivalence of the two integrals.
Considering the above observation, to prove that the BDS integral is a special case of a type of the net Riemann integral, we shall use the following simplification of the decomposition of (strongly) measurable functions given in (Musiał, 1991, Lemma 5.1, p. 201):
Proposition 5.6.
If $f:\Omega\rightarrow\mathbb{R}$ is a Borel measurable function, then there exits a bounded (Borel) measurable function $g:\Omega\rightarrow\mathbb{R}$ and a(n) (elementary and Borel) measurable function $h:\Omega\rightarrow\mathbb{R}$ of the form $\sum_{i=1}^{\infty}x_{n}\mathbbm{1}_{A_{i}}$, with pairwise disjoint $A_{i}\in\mathcal{H}$ such that:
$f=g+h.$
As is well known that each bounded real $\mu$-measurable function is a uniform limit of simple functions defined on $\mathcal{H}$ (see chapter 1 of Dinculeanu (2000)) and that $\mu$ is of finite semivariation by Proposition 5.5, we get immediately by Theorem 4.12, Observation 5.1 and the fact that the $S^{*}$-partition integral equals de BDS-integral on simple functions that:
Proposition 5.7.
Let $g:\Omega\rightarrow\mathbb{R}$ a bounded $\mu$-measurable function. Then, $g$ is $S^{*}$-partition integrable and its integral equals the $BDS$-integral of gg [^22] Which is always BDS-integrable by construction..
Now, we need the next result which is contained in (Drewnowski and Labuda, 2013, Remarks 3.3. (3), p. 627):
Proposition 5.8.
Let $h:\Omega\rightarrow\mathbb{R}$ be a(n) (elementary) $\mu$-measurable function of the form $\sum_{i=1}^{\infty}x_{n}\mathbbm{1}_{A_{i}}$, with pairwise disjoint $A_{i}\in\mathcal{H}$ which is BDS-integrable. Then,
$\int_{A}^{(BDS)}hd\mu=\sum_{i=1}^{\infty}x_{i}\mu(A\cap A_{i}),$
for each $A\in\mathcal{H}$ where the series is unconditionally summable in $X$.
Combining these results, we get:
Theorem 5.4 (BDS integrability implies $S^{*}$-partition integrability).
Let $f\in L^{0}(\mu)$ be a BDS-integrable function. Then, $f$ is $S^{*}$-partition integrable and:
$\int_{A}^{S^{*}}fd\mu=\int_{A}^{(BDS)}fd\mu,$
for each $A\in\mathcal{H}$.
Proof.
As $f$ is $\mu$-measurable, by Observation 5.1, we may assume that $f$ is (Borel) measurable. Therefore, by Proposition 5.6, there exits a bounded measurable function $g:\Omega\rightarrow\mathbb{R}$ and a(n) (elementary) measurable function $h:\Omega\rightarrow\mathbb{R}$ of the form $\sum_{i=1}^{\infty}x_{n}\mathbbm{1}_{A_{i}}$, with pairwise disjoint $A_{i}\in\mathcal{H}$ such that:
$f=g+h,$
As $f$ is BDS-integrable and $g$ is bounded, and therefore $BDS$-integrable by Proposition 5.7 with BDS-integral equal to its $S^{*}$-partition integral, it follows that $h$ is BDS-integrable and, by Proposition 5.8,
$\int_{A}^{(BDS)}hd\mu=\sum_{i=1}^{\infty}x_{i}\mu(A\cap A_{i}),$
for $A\in\mathcal{H}$ and with the right-hand series unconditionally summable in $X$.
By the same techniques as in the Pavlakos integral of elementary functions in Proposition 5.1, we get, using the unconditional summability of the above series, that $h$ is $S^{*}$-partition integrable and:
$\int_{A}^{S^{*}}hd\mu=\sum_{i=1}^{\infty}x_{i}\mu(A\cap A_{i}).$
Combining with the results on $g$ and using the additivity of each integral, we get the result. ∎
Therefore, by Proposition 3.8, we get that the BDS-integral is a special case of the $S^{*}$-partition and Sion integrals, and therefore a special case of the Net Riemann integral. All the results above can be developed for $\delta$-rings in the place of $\sigma$-algebras, as detailed in Thomas (2012) for the BDS-integral with the use of Theorem 4.13 accounting for the $\delta$-ring. One consequence is that the (stochastic) integral of Schwartz (1981), as well as the integrals of Masani and Niemi (1989a), Masani and Niemi (1989b) and Masani and Niemi (1992), are contained as special cases of the Riemmanian definition of integration given by the net Riemann integral.
We now study the next point in relation to Haluška (1988) bornological integral.
5.2.2 The Haluska and Rodriguez-Salazar integrals in Bornological Locally Convex Spaces.
For this section, we mainly follow the article Duchoň and Haluška (2005) for the basic elements in establishing the bornological integral.
In this case, we let $X,Y$ be two complete bornological locally convex topological vector spaces with respect to the field of real numbers, with bornologies $\mathcal{B}_{X}$ and $\mathcal{B}_{Y}$. Let also be $\mathcal{U}$ and $\mathcal{W}$ the set of all Banach disks in $X$ and $Y$ respectively, in such a way that $X$ and $Y$ are inductive limits of (separable) Banach spaces $X_{U}$ and $Y_{W}$,
$X=inj\lim_{U\in\mathcal{U}}X_{U},$
$Y=inj\lim_{W\in\mathcal{W}}Y_{W},$
We equip both spaces with the net convergence structure of Mackey convergence on each Banach disk (see O’Brien et al. (2023) for details of the net convergence structures associated to bornologies).
We also denote $p_{U}:X\rightarrow[0,\infty]$ the Minkowski functional associated to each member $U\in\mathcal{U}$ (similarly for $\mathcal{W}$), and by $L(X,Y)$ the space of all continuous linear operators between $X$ and $Y$.
Further, for this section, let $\mathcal{H}$ be a $\delta$-ring of subsets of $\Omega$ and $\sigma(\mathcal{H})$ the $\sigma$-algebra generated by $\mathcal{H}$.
In this case, we may define a central notion for the theory of bilinear integral of Dobrakov type:
Definition 5.6 (Bilinear Semivariation and Sets of Finite Semivariation).
Let $\mu:\mathcal{H}\rightarrow L(X,Y)$ be a finitely additive operator valued measure. For $(U,W)\in\mathcal{U}\times\mathcal{W}$, we denote by $\hat{\mu}_{U,W}$ the $(U,W)$-semivariation of $\mu$, that is, the order continuous (extended real) submeasure (defined on $\mathcal{H}$) given by:
$\hat{\mu}_{U,W}(A)=\sup\{p_{W}(\sum_{i=1}^{n}\mu(A\cap A_{i})x_{i}):\{x_{i}\}_{i=1}^{n}\in U;\{A_{i}\}_{i=1}^{n}\ \text{disjoint in}\ \mathcal{H},n\in\mathbb{N}\}.$
We then denote by $\mathcal{H}_{U,W}$ the largest $\delta$-ring of sets from $\mathcal{H}$ which are of finite $\hat{\mu}_{U,W}$ semivariation. Also, $\hat{\mu}_{\mathcal{U},\mathcal{W}}$ will be the set of all semivariations $\hat{\mu}_{U,W}$.
We now need some analogous constructions ,in the bornological case, of operator valued measures that are $\sigma$-additive in the strong operator topology of a Banach space. For that, we need some further notions related to operators between bornological spaces.
In this context, let $\Phi$ be the class of all functions defined on $\mathcal{U}$, and with values in $\mathcal{W}$, equipped with the partial order given by: for $\phi,\zeta\in\Phi$, we say that $\phi\leq\zeta$ if and only if $\phi(U)\subseteq\zeta(U)$ for each $U\in\mathcal{U}$. We also suppose that $\Phi\subseteq L(X,Y)$.
We then have the next measure related structures from Duchoň and Haluška (2005):
Definition 5.7 (Countable Additivity and Bornologies).
For $\mu:\mathcal{H}\rightarrow L(X,Y)$ a finitely additive operator valued measure, we say that:
@@
    For $(U,W)\in\mathcal{U}\times\mathcal{W}$, $\mu$ is of $\sigma$-finite $(U,W)$-semivariation if there exists a sequence $\{A_{n}\}_{n\in\mathbb{N}}$ in $\mathcal{H}_{(U,W)}$ such that $\Omega=\bigcup_{n=1}^{\infty}A_{n}$.
@@
    For $\phi\in\Phi$, $\mu$ is of $\sigma_{\phi}$-finite $(\mathcal{U},\mathcal{W})$-semivariation if for every $U\in\mathcal{U}$, $\mu$ is of $\sigma$-finite $(\mathcal{U},\phi(\mathcal{U}))$-semivariation.